don t know if you can help Im looking for a chart that will give me a break down of measurments from 35 15/16 to 65 15/16 in half for example
52 11/16 = 26 3/8 Im looking for a printable chart can you help or tell me were to look

Answer by addingup(3677)

Not sure what you're after, but one-half of 52 11/16 is 26 11/32 and not 26 3/8:
52 11/16*1/2 = 843/16*1/2 = 26 11/32
:
If your calculator doesn't do fractions you can use an online calculator like this one: http://www.calculatorsoup.com/calculators/math/fractions.php